More Vectors! Let A = 1.2 cmi + 5.4cmj. B-1.3 emi + 2.4 cm3- : 4.4 cmi Find: a. The magnitude and direction of A,B, and C. Express the direction as an angle, in degrees, assuming 0 is the positive axis and counterclockwise angles are positive. A: Magnitude B: Magnitude = C: Magnitude cm Direction cm Direction = i cm Direction b. The magnitude and direction of the vector R-A 1.5B+3C. Magnitude R i cm Direction
Answer #1

A = ( 1.2i + 5.4j) Cmm

A=sqrt{A_{x}^{2}+A_{y}^{2}}

A = V( 1.2 )2 + (5.4)2

ANSWER: 5.532cm

Direction θ tan-1 (Ay/Ar )

0tan(5.4cm/1.2cm)

ANSWER: θ = 77 .47

=======================

vec{B}=(1.3i+2.4j)cm

B=sqrt{B_{x}^{2}+B_{y}^{2}}

B=sqrt{(1.3)^{2}+(2.4)^{2}}

ANSWER: {color{Red} B=2.729cm}

Direction heta =tan^{-1}(B_{y}/B_{x})

heta =tan^{-1}(2.4cm/1.3cm)

ANSWER: θ = 61.56

=======================

C - (4.4i)cm

C=sqrt{C_{x}^{2}+C_{y}^{2}}

C = V/(4.4)2 (O)2

ANSWER: {color{Red} C=4.4cm}

Direction θ = tan-1 (Cyc,)

heta =tan^{-1}(0/4.4cm)

ANSWER: {color{Red} heta =0^{circ}}

=======================

vec{R}=vec{A}-1.5vec{B}+3vec{C}

vec{R}=(1.2i+5.4j)-1.5(1.3i+2.4j)+3(4.4i)

vec{R}=(1.2i+5.4j)-(1.95i+3.6j)+(13.2i)

R = 1.2i 5.4) _ 1.951-3.6 j 13.2ǐ

vec{R}=(1.2-1.95+13.2)i+(5.4-3.6)j

R12.45i + 1.8

2

R=sqrt{(12.45)^{2}+(1.8)^{2}}

ANSWER: {color{Red} R=12.579cm}

Direction heta =tan^{-1}(R_{y}/R_{x})

heta =tan^{-1}(1.8cm/12.45cm)

ANSWER: {color{Red} heta =8.227^{circ}}

=======================